# Stochastic Cahn–Hilliard Equations from One-Dimensional Ising–Kac–Kawasaki Dynamics

### 31 Mar 2026 v1

@@ 4.2 Two-block replacement for the conserved field
$\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(\widetilde{m}_{x}^{\ell,L}(\sigma(t))-m_{x}^{L}(\sigma(t))\Bigr)\,dt\right|\leq CT\|J\|_{\infty}\frac{\ell}{L}.$
Since $\ell/L\to 0$, the result follows. ∎
Corollary 7 (Comparison of local equilibrium averages at two block scales).
Assume, in addition, that the family $\Phi_{\Psi,\ell}^{\gamma}$ satisfies:
@@
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(\Phi_{\Psi,\ell_{\gamma}}^{\gamma}(m_{x}^{\ell_{\gamma}}(\sigma(t)))-\Phi_{\Psi,L_{\gamma}}^{\gamma}(m_{x}^{L_{\gamma}}(\sigma(t)))\Bigr)\,dt\right|=0.$
@@
$\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{\ell})-\Phi_{\Psi,L}^{\gamma}(m_{x}^{L})=\Bigl[\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{\ell})-\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{L})\Bigr]+\Bigl[\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{L})-F_{\Psi}^{\gamma}(m_{x}^{L})\Bigr]+\Bigl[F_{\Psi}^{\gamma}(m_{x}^{L})-\Phi_{\Psi,L}^{\gamma}(m_{x}^{L})\Bigr].$
@@ 4.3 Kac-scale regularity of mesoscopic block averages
In particular, if $\gamma^{2}L^{3}\to 0$, then $\Phi_{\Psi,L}^{\gamma}(m)=a_{0,L,\gamma}+a_{2,L,\gamma}m^{2}+\rho_{\gamma,L}(m),$ with
@@
Under $\nu_{L,m}^{\mathrm{aux}}$, exchangeability implies that for every $i\neq j$, $\mathbb{E}_{\nu_{L,m}^{\mathrm{aux}}}[\eta_{i}\eta_{j}]=:c_{L}(m).$ Now
@@
We now compare $\nu_{L,m}^{\mathrm{aux}}$ with the true Ising–Kac local canonical Gibbs measure. Write
$H_{\gamma,L}^{\mathrm{blk}}(\eta)=-\frac{1}{2}\sum_{u,v\in\Lambda_{L}}\kappa_{\gamma}(u-v)\eta_{u}\eta_{v}.$
Set $\bar{\kappa}_{\gamma,L}:=\gamma\,\mathfrak{K}(0).$ Then
$H_{\gamma,L}^{\mathrm{blk}}(\eta)=-\frac{1}{2}\bar{\kappa}_{\gamma,L}\left(\sum_{u\in\Lambda_{L}}\eta_{u}\right)^{2}-\frac{1}{2}\sum_{u,v\in\Lambda_{L}}\bigl(\kappa_{\gamma}(u-v)-\bar{\kappa}_{\gamma,L}\bigr)\eta_{u}\eta_{v}.$
On the fixed-magnetization sector
$\Omega_{L,m}:=\left\{\eta\in\{-1,1\}^{\Lambda_{L}}:\ \sum_{u\in\Lambda_{L}}\eta_{u}=m(2L+1)\right\},$
the first term is constant. Hence the configuration dependence is entirely contained in
$R_{\gamma,L}(\eta):=-\frac{1}{2}\sum_{u,v\in\Lambda_{L}}\bigl(\kappa_{\gamma}(u-v)-\bar{\kappa}_{\gamma,L}\bigr)\eta_{u}\eta_{v}.$
@@
Lemma 11 (Kac-scale regularity of mesoscopic block averages).
@@
Then, for every $J\in C^{\infty}(\mathbb{T})$,
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}\left|J\!\left(\frac{x}{N}\right)\right|\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(x-z)\bigl(m_{z}^{L_{\gamma}}(\sigma(t))-\bar{h}_{\gamma,L_{\gamma}}(\sigma(t);x)\bigr)^{2}\,dt\right]=0.$
@@
For brevity write
$m_{z}^{L}(t):=m_{z}^{L_{\gamma}}(\sigma(t)),\qquad\bar{h}_{\gamma,L}(t;x):=\bar{h}_{\gamma,L_{\gamma}}(\sigma(t);x).$
Set
$\mathcal{R}_{\gamma}:=\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}\left|J\!\left(\frac{x}{N}\right)\right|\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(x-z)\bigl(m_{z}^{L}(t)-\bar{h}_{\gamma,L}(t;x)\bigr)^{2}\,dt.$
Since $\kappa_{\gamma}\geq 0$ and $\sum_{z}\kappa_{\gamma}(x-z)=1$, $\bar{h}_{\gamma,L}(t;x)$ is the $\kappa_{\gamma}(x-\cdot)$-weighted mean of $\{m_{z}^{L}(t)\}_{z}$. Hence, by the minimizing property of the weighted mean,
$\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(x-z)\bigl(m_{z}^{L}(t)-\bar{h}_{\gamma,L}(t;x)\bigr)^{2}\leq\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(x-z)\bigl(m_{z}^{L}(t)-m_{x}^{L}(t)\bigr)^{2}.$
Since $|m_{z}^{L}(t)|\leq 1$,
$\bigl(m_{z}^{L}(t)-m_{x}^{L}(t)\bigr)^{2}\leq 2\,|m_{z}^{L}(t)-m_{x}^{L}(t)|.$
Therefore
$\mathcal{R}_{\gamma}\leq 2\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}\left|J\!\left(\frac{x}{N}\right)\right|\sum_{y\in\Lambda_{N}}\kappa_{\gamma}(y)\,|m_{x+y}^{L}(t)-m_{x}^{L}(t)|\,dt.$
Fix $A>0$, and define $M_{\gamma}^{A}:=\left\lfloor\frac{A}{\gamma}\right\rfloor.$ Split the Kac sum as
$\sum_{y}\kappa_{\gamma}(y)|m_{x+y}^{L}-m_{x}^{L}|=\sum_{|y|\leq M_{\gamma}^{A}}\kappa_{\gamma}(y)|m_{x+y}^{L}-m_{x}^{L}|+\sum_{|y|>M_{\gamma}^{A}}\kappa_{\gamma}(y)|m_{x+y}^{L}-m_{x}^{L}|.$
and
$\mathcal{R}_{\gamma}\leq 2\mathcal{R}_{\gamma,A}^{\mathrm{near}}+2\mathcal{R}_{\gamma,A}^{\mathrm{tail}}.$
For the near-field part, since $\mathfrak{K}\in L^{\infty}$, $\sup_{y}\kappa_{\gamma}(y)\leq\gamma\|\mathfrak{K}\|_{L^{\infty}}.$ Hence
$\mathcal{R}_{\gamma,A}^{\mathrm{near}}\leq C_{A}\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}\frac{1}{2M_{\gamma}^{A}+1}\sum_{|y|\leq M_{\gamma}^{A}}\left|J\!\left(\frac{x}{N}\right)\right||m_{x+y}^{L_{\gamma}}(\sigma(t))-m_{x}^{L_{\gamma}}(\sigma(t))|\,dt.$
Since
$M_{\gamma}^{A}\to\infty,\qquad\frac{L_{\gamma}}{M_{\gamma}^{A}}\sim\frac{\gamma L_{\gamma}}{A}\to 0,\qquad\frac{M_{\gamma}^{A}}{N}\sim\frac{A}{\gamma N}\to 0,$
Theorem 6 applies with small block size $L_{\gamma}$ and larger averaging window $M_{\gamma}^{A}$, so for every fixed $A>0$,
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}[\mathcal{R}_{\gamma,A}^{\mathrm{near}}]=0.$
For the tail part, since $|m_{x+y}^{L}-m_{x}^{L}|\leq 2$,
$\mathcal{R}_{\gamma,A}^{\mathrm{tail}}\leq 2T\|J\|_{\infty}\sum_{|y|>M_{\gamma}^{A}}\kappa_{\gamma}(y).$
Using $\kappa_{\gamma}(y)=\gamma\mathfrak{K}(\gamma y)$ and $\mathfrak{K}\in L^{1}$,
$\limsup_{\gamma\downarrow 0}\sum_{|y|>M_{\gamma}^{A}}\kappa_{\gamma}(y)\leq C\int_{|u|>A/2}|\mathfrak{K}(u)|\,du,$
@@
We need the following basic estimates for $X_{\gamma}$.
Lemma 12.
Under scale assumption, the discrete random field $X_{\gamma}$ satisfies
$\left\|\nabla_{\varepsilon_{\gamma}}X_{\gamma}\right\|_{L^{\infty}}\leq C\,\frac{\gamma}{\delta_{\gamma}\varepsilon_{\gamma}},$ (27)
and
$\sup_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\left\|X_{\gamma}\right\|_{L^{p}}dt\right]<\infty.$ (28)
Proof.
Since $X_{\gamma}(t,\varepsilon_{\gamma}i)=\delta_{\gamma}^{-1}\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)\sigma_{j}\big(\frac{t}{\alpha_{\gamma}}\big),$ we have
$\left|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\right|=\frac{1}{\delta_{\gamma}\varepsilon_{\gamma}}\left|h_{\gamma}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i+1\Big)-h_{\gamma}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i\Big)\right|.$
By $\kappa_{\gamma}(z)=\gamma\,\mathfrak{K}(\gamma z),\mathfrak{K}\in W^{1,\infty}(\mathbb{R}),$ and $|\sigma(j)|\leq 1$, we obtain
$\left|h_{\gamma}(\sigma;i+1)-h_{\gamma}(\sigma;i)\right|\leq\sum_{j\in\Lambda_{N}}\bigl|\kappa_{\gamma}(i+1-j)-\kappa_{\gamma}(i-j)\bigr|\leq C\,\gamma.$
Therefore
$\sup_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\Big(\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}\Big|X_{\gamma}(t,\varepsilon_{\gamma}i)\Big|^{p}\Big)^{\frac{1}{p}}dt\right]\leq\sup_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\sup_{i\in\Lambda_{N}}\Big|X_{\gamma}(t,\varepsilon_{\gamma}i)\Big|dt\right]<\infty.$
this becomes estimate (28) ∎
Now we derive the nonlinear drift for $\Psi=d_{0}^{2}$ from the the second-order Boltzmann–Gibbs principle.
@@
By Theorem 5, Corollary 7, and Lemma 10, the local observable $d_{i}^{2}$ can be replaced by the mesoscopic local equilibrium average By Lemma 10,
$\Phi_{\Psi,L_{\gamma}}^{\gamma}\bigl(m_{i}^{L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big)\Big)\bigr)=a_{0,\gamma}+a_{2,\gamma}\big(m_{i}^{L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big)\Big)\big)2+\rho_{\gamma,L_{\gamma}}(m_{i}^{L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big)\Big)),$
with
@@
$\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi;t)=-\frac{\beta}{4}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Bigl(a_{0,\gamma}+a_{2,\gamma}(m_{i}^{L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big)\Big))^{2}\Bigr)\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)+r_{\gamma}^{(1)}(\phi;t)+r_{\gamma}^{(2)}(\phi;t),$
@@
$r_{\gamma}^{(2)}(\phi;t):=-\frac{\beta}{4}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\rho_{\gamma,L_{\gamma}}\bigl(m_{i}^{L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big)\Big)\bigr)\,\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i).$
@@
Combining with $|m_{i}^{L_{\gamma}}|,|\bar{h}_{\gamma,L_{\gamma}}(i)|,|h_{\gamma}(i)|\leq 1$, we have
$(m_{i}^{L_{\gamma}})^{2}-h_{\gamma}(i)^{2}=$ $\bigl((m_{i}^{L_{\gamma}})^{2}-\bar{h}_{\gamma,L_{\gamma}}(i)^{2}\bigr)+\bigl(\bar{h}_{\gamma,L_{\gamma}}(i)^{2}-h_{\gamma}(i)^{2}\bigr)$
$\leq$ $2|m_{i}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)|+2|\bar{h}_{\gamma,L_{\gamma}}(i)-h_{\gamma}(i)|.$
@@
$\left|\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i}\Bigl[(m_{i}^{L_{\gamma}})^{2}-h_{\gamma}(i)^{2}\Bigr]\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)dt\right|\leq 2\mathcal{E}_{\gamma}^{(1)}+2\mathcal{E}_{\gamma}^{(2)},$
@@
$\mathcal{E}_{\gamma}^{(1)}:=\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}|m_{i}^{L_{\gamma}}(\sigma\big(\frac{t}{\alpha_{\gamma}}\big))-\bar{h}_{\gamma,L_{\gamma}}(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i)|\,|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|\,dt,$
@@
$\mathcal{E}_{\gamma}^{(2)}:=\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Big|\bar{h}_{\gamma,L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i\Big)-h_{\gamma}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i\Big)\Big||\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|\,dt.$

### 05 Apr 2026 v2

@@ 4.2 Two-block replacement for the conserved field
$\left|\int_{0}^{T}\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}J\left(\varepsilon_{\gamma}i\right)\Bigl|\widetilde{m}_{i}^{\ell,L}(\sigma^{N}(\alpha_{\gamma}^{-1}t))-m_{i}^{L}(\sigma^{N}(\alpha_{\gamma}^{-1}t))\Bigr|dt\right|\leq CT\|J\|_{\infty}\,\frac{\ell}{L}.$
Since $\ell_{\gamma}/L_{\gamma}\to 0$, we obtain
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}J\left(\varepsilon_{\gamma}i\right)\Bigl|\widetilde{m}_{i}^{\ell_{\gamma},L_{\gamma}}(\sigma^{N}(\alpha_{\gamma}^{-1}t))-m_{i}^{L_{\gamma}}(\sigma^{N}(\alpha_{\gamma}^{-1}t))\Bigr|dt\right]=0.$ (32)
Combining (31) and (32), we conclude the estimate (29). ∎
We compare the local equilibrium averages at two block scales.
Corollary 7.
Assume that the family $\Phi_{\Psi,\ell}^{\gamma}$ satisfies:
@@
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}J\left(\varepsilon_{\gamma}i\right)\Bigl(\Phi_{\Psi,\ell_{\gamma}}^{\gamma}(m_{i}^{\ell_{\gamma}}(\sigma^{N}(t)))-\Phi_{\Psi,L_{\gamma}}^{\gamma}(m_{i}^{L_{\gamma}}(\sigma^{N}(t)))\Bigr)\,dt\right|=0.$
@@
$\Phi_{\Psi,\ell}^{\gamma}(m_{i}^{\ell})-\Phi_{\Psi,L}^{\gamma}(m_{i}^{L})=\Bigl[\Phi_{\Psi,\ell}^{\gamma}(m_{i}^{\ell})-\Phi_{\Psi,\ell}^{\gamma}(m_{i}^{L})\Bigr]+\Bigl[\Phi_{\Psi,\ell}^{\gamma}(m_{i}^{L})-F_{\Psi}^{\gamma}(m_{i}^{L})\Bigr]+\Bigl[F_{\Psi}^{\gamma}(m_{i}^{L})-\Phi_{\Psi,L}^{\gamma}(m_{i}^{L})\Bigr].$
@@ 4.3 Kac-scale regularity of mesoscopic block averages
In particular, if $\gamma^{2}L^{3}\to 0$, then
$\Phi_{\Psi,L}^{\gamma}(m)=a_{0,L,\gamma}+a_{2,L,\gamma}m^{2}+\rho_{\gamma,L}(m),$ (33)
with
@@
Under the auxiliary canonical measure $\nu_{L,m}^{\mathrm{aux}}$, exchangeability implies that for every $i\neq j$, $\mathbb{E}_{\nu_{L,m}^{\mathrm{aux}}}[\eta_{i}\eta_{j}]=:c_{L}(m).$ Now
@@
We now compare $\nu_{L,m}^{\mathrm{aux}}$ with the original local canonical Gibbs measure $\mu_{\gamma,L,m}^{\mathrm{can}}$. We rewrite the block Hamiltonian
$H_{\gamma,L}^{\mathrm{blk}}(\eta)=-\frac{1}{2}\sum_{u,v\in\Lambda_{L}}\kappa_{\gamma}(u-v)\eta_{u}\eta_{v}=-\frac{1}{2}\gamma\,\mathfrak{K}(0)\left(\sum_{u\in\Lambda_{L}}\eta_{u}\right)^{2}-\frac{1}{2}\sum_{u,v\in\Lambda_{L}}\bigl(\kappa_{\gamma}(u-v)-\gamma\,\mathfrak{K}(0)\bigr)\eta_{u}\eta_{v}.$
On the fixed-magnetization sector $\Omega_{L,m}:=\left\{\eta\in\{-1,1\}^{\Lambda_{L}}:\ \sum_{u\in\Lambda_{L}}\eta_{u}=m(2L+1)\right\},$ the first term is constant. Hence the configuration dependence is entirely contained in
$R_{\gamma,L}(\eta):=-\frac{1}{2}\sum_{u,v\in\Lambda_{L}}\bigl(\kappa_{\gamma}(u-v)-\gamma\,\mathfrak{K}(0)\bigr)\eta_{u}\eta_{v}.$
@@
Now we consider the Kac-scale regularity of mesoscopic block averages.
Lemma 13.
For every $J\in L^{\infty}(\mathbb{T})$, we have
$\lim_{\gamma\downarrow 0}\left|\mathbb{E}_{f_{N,0}}\int_{0}^{T}\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}J\!\left(\varepsilon_{\gamma}i\right)\Big[\Big(m_{i}^{L_{\gamma}}\big(\sigma^{N}\big(t/\alpha_{\gamma}\big)\big)\Big)^{2}-\Big(\bar{h}_{\gamma,L_{\gamma}}(\sigma^{N}\big(t/\alpha_{\gamma}\big);i)\Big)^{2}\Big]\,dt\right|=0.$
@@
For brevity, we write
$m_{z}^{L}(t):=m_{z}^{L_{\gamma}}(\sigma(t/\alpha_{\gamma})),\qquad\bar{h}_{\gamma,L}(t;x):=\bar{h}_{\gamma,L_{\gamma}}(\sigma(t/\alpha_{\gamma});x)=\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(x-z)m_{z}^{L_{\gamma}}(t/\alpha_{\gamma}).$
Since $\kappa_{\gamma}\geq 0$ and $\sum_{z}\kappa_{\gamma}(x-z)=1$, $\bar{h}_{\gamma,L}(t;x)$ is the $\kappa_{\gamma}(x-\cdot)$-weighted mean of $\{m_{z}^{L}(t)\}_{z}$. Then
$\mathcal{R}_{\gamma}:=$ $\left|\int_{0}^{T}\varepsilon_{\gamma}\sum_{x\in\Lambda_{N}}J\!\left(\varepsilon_{\gamma}x\right)\Big[\bigl(m_{x}^{L}(t)\bigr)^{2}-\bigl(\bar{h}_{\gamma,L}(t;x)\bigr)^{2}\Big]dt\right|$
$=$ $\left|\int_{0}^{T}\varepsilon_{\gamma}\sum_{x\in\Lambda_{N}}J\!\left(\varepsilon_{\gamma}x\right)\bigl(m_{x}^{L}(t)+\bar{h}_{\gamma,L}(t;x)\bigr)\bigl(m_{x}^{L}(t)-\bar{h}_{\gamma,L}(t;x)\bigr)dt\right|$
$=$ $\left|\int_{0}^{T}\varepsilon_{\gamma}\sum_{x\in\Lambda_{N}}J\!\left(\varepsilon_{\gamma}x\right)\bigl(m_{x}^{L}(t)+\bar{h}_{\gamma,L}(t;x)\bigr)\sum_{y\in\Lambda_{N}}\kappa_{\gamma}(y)\bigl(m_{x+y}^{L}(t)-m_{x}^{L}(t)\bigr)dt\right|$ (34)
For each fix $A>0$, we define $M_{\gamma}^{A}:=\left\lfloor\frac{A}{\gamma}\right\rfloor.$ We split the Kac sum as
$\sum_{y}\kappa_{\gamma}(y)(m_{x+y}^{L}-m_{x}^{L})=\sum_{|y|\leq M_{\gamma}^{A}}\kappa_{\gamma}(y)(m_{x+y}^{L}-m_{x}^{L})+\sum_{|y|>M_{\gamma}^{A}}\kappa_{\gamma}(y)(m_{x+y}^{L}-m_{x}^{L}),$
so that $\mathcal{R}_{\gamma}=\mathcal{R}_{\gamma,A}^{\mathrm{near}}+\mathcal{R}_{\gamma,A}^{\mathrm{tail}}$.
For the near-field part $\mathcal{R}_{\gamma,A}^{\mathrm{near}}$, since $\mathfrak{K}\in L^{\infty}$, we have $\sup_{y}\kappa_{\gamma}(y)\leq\gamma\|\mathfrak{K}\|_{L^{\infty}}.$ By estimate (31) in Theorem 6 applies with small block size $L_{\gamma}$ and larger averaging window $M_{\gamma}^{A}$, for every fixed $A>0$ we have
$\mathcal{R}_{\gamma,A}^{\mathrm{near}}=$ $C_{A}\Big|\int_{0}^{T}\varepsilon_{\gamma}\sum_{x\in\Lambda_{N}}J\!\left(\varepsilon_{\gamma}x\right)\bigl(m_{x}^{L}(t)+\bar{h}_{\gamma,L}(t;x)\bigr)\sum_{|y|\leq M_{\gamma}^{A}}\kappa_{\gamma}(y)\Big(m_{x+y}^{L}(\sigma(t))-m_{x}^{L}(\sigma(t))\Big)dt\Big|$
$\leq$ $C\|J(m^{L}+\bar{h}_{\gamma,L})\|_{L^{\infty}}A\sqrt{T\alpha_{\gamma}A/\gamma}$ (35)
Fix $A>0$ and let $\gamma\to 0$, so that $A^{\frac{3}{2}}\big(\frac{\alpha_{\gamma}}{\gamma}\big)^{\frac{1}{2}}\to 0$ and Since $\|m^{L}+\bar{h}_{\gamma,L}\|_{L^{\infty}}\leq 2$, we have $\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}[\mathcal{R}_{\gamma,A}^{\mathrm{near}}]=0.$
For the tail part, since $|m_{x+y}^{L}-m_{x}^{L}|\leq 2$, $\kappa_{\gamma}(y)=\gamma\mathfrak{K}(\gamma y)$, and $\mathfrak{K}\in L^{1}$, we have
$\mathcal{R}_{\gamma,A}^{\mathrm{tail}}=$ $C\Big|\int_{0}^{T}\varepsilon_{\gamma}\sum_{x\in\Lambda_{N}}J\!\left(\varepsilon_{\gamma}x\right)\bigl(m_{x}^{L}(t)+\bar{h}_{\gamma,L}(t;x)\bigr)\sum_{|y|>M_{\gamma}^{A}}\kappa_{\gamma}(y)\Big(m_{x+y}^{L}(\sigma(t))-m_{x}^{L}(\sigma(t))\Big)dt\Big|$
$\leq$ $C\|J\|_{\infty}\sum_{|y|>M_{\gamma}^{A}}\kappa_{\gamma}(y)$
$\leq$ $C\|J\|_{\infty}\int_{|u|>A/2}|\mathfrak{K}(u)|\,du,$
@@
Now we derive the nonlinear drift for $\Psi=d_{0}^{2}$ from the second-order Boltzmann–Gibbs principle.
@@
By Theorem 5, Corollary 7, and Lemma 12, the local observable $d_{i}^{2}$ can be replaced by the mesoscopic local equilibrium average
$\Phi_{\Psi,L_{\gamma}}^{\gamma}\bigl(m_{i}^{L_{\gamma}}\Big(\sigma^{N}\big(\frac{t}{\alpha_{\gamma}}\big)\Big)\bigr)=a_{0,\gamma}+a_{2,\gamma}\Big(m_{i}^{L_{\gamma}}\Big(\sigma^{N}\big(\frac{t}{\alpha_{\gamma}}\big)\Big)\Big)^{2}+\rho_{\gamma,L_{\gamma}}(m_{i}^{L_{\gamma}}\Big(\sigma^{N}\big(\frac{t}{\alpha_{\gamma}}\big)\Big)),$
where
@@
$\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi;t)=-\frac{\beta}{4}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Bigl[a_{0,\gamma}+a_{2,\gamma}\Big(m_{i}^{L_{\gamma}}\big(\sigma(t/\alpha_{\gamma})\big)\Big)^{2}\Bigr]\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)+r_{\gamma}^{(1)}(\phi;t)+r_{\gamma}^{(2)}(\phi;t),$
@@
$r_{\gamma}^{(2)}(\phi;t):=-\frac{\beta}{4}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\rho_{\gamma,L_{\gamma}}\bigl(m_{i}^{L_{\gamma}}\Big(\sigma^{N}\big(\frac{t}{\alpha_{\gamma}}\big)\Big)\bigr)\,\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i).$
@@
$\left|\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i}\Bigl[\Big(m_{i}^{L_{\gamma}}\big(\sigma^{N}\big(t/\alpha_{\gamma}\big)\big)\Big)^{2}-\Big(h_{\gamma}(\sigma^{N}\big(t/\alpha_{\gamma}\big);i)\Big)^{2}\Bigr]\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)dt\right|\leq 2\mathcal{E}_{\gamma}^{(1)}+2\mathcal{E}_{\gamma}^{(2)},$
@@
$\mathcal{E}_{\gamma}^{(1)}:=\left|\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Big[\Big(m_{i}^{L_{\gamma}}\big(\sigma^{N}\big(t/\alpha_{\gamma}\big)\big)\Big)^{2}-\Big(\bar{h}_{\gamma,L_{\gamma}}(\sigma^{N}\big(t/\alpha_{\gamma}\big);i)\Big)^{2}\Big]\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)dt\right|,$
@@
$\mathcal{E}_{\gamma}^{(2)}:=\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Big|\Big(\bar{h}_{\gamma,L_{\gamma}}(\sigma^{N}\big(t/\alpha_{\gamma}\big);i)\Big)^{2}-\Big(h_{\gamma}(\sigma^{N}\big(t/\alpha_{\gamma}\big);i)\Big)^{2}\Big||\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|\,dt.$